OpenStudy (anonymous):

what is the sum of the first 10 terms of the sequence defined by an=2n-3

OpenStudy (shamil98):

To find the sum s(10) you must first find what 2(1)-3 and 2(10)-3 equals. so it is -1 and 17. s(10) = 1/2(-1+17)(10) s(10) = 8(10) s(10 = 80